OpenStudy (anonymous):

How can I condense the expression? 1/2Log(x+5)-2Logx

OpenStudy (anonymous):

1/2Log(x+5)-2Log(x) =Log(x+5)^(1/2)-Log(x^2) =Log(sqrt(x+5)/x^2)